OpenStudy (anonymous):

I've searched the internet with no success. Please help. If f(x) = x^2 + 1 and g(x) = 3x + 1, find [f(2)-g(1)]^2

OpenStudy (solomonzelman):

first find f(2) - plug in 2 for x into the x^2+1

OpenStudy (solomonzelman):

then find g(1) - plug in 1 for x into the 3x+1

OpenStudy (solomonzelman):

tell me what you get for f(2) and g(1) (respectively) please...

OpenStudy (anonymous):

So f(2) would plug in and make it 2^2 +1 = 5 and g(1) would plug in to make it 3(1) + 1 = 4, but how would the power at the end work?

OpenStudy (solomonzelman):

f(2)=5 g(1)=4 so then [ f(2)-g(1) ] ^2 becomes [5 - 4]^2

OpenStudy (solomonzelman):

and that is equal t☼ ?

OpenStudy (anonymous):

25 - 16 = 9

OpenStudy (anonymous):

Thanks a thousand!

OpenStudy (solomonzelman):

no no

OpenStudy (solomonzelman):

you are first subtracting f(2)-g(1) (that is subtract 5-4) AND only then take the second power

OpenStudy (solomonzelman):

(A-B)^2 is not same as A^2-B^2

OpenStudy (anonymous):

So then you would have 1^2?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and that is same as ?

OpenStudy (anonymous):

One.

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

Okay. Thanks for catching me there.

OpenStudy (solomonzelman):

sure, not a problem!

OpenStudy (phi):

to do the arithmetic correctly, remember *order of operations* people learn various mnemonics PEDMAS for example, which means you do the first on the list before moving on to the next: parens, exponents, divide/multiply, add/subtract thus with (5-4)^2 you do the stuff inside the parens before doing the exponents 1^2 now you can do the exponent: 1*1 = 1